OpenStudy (anonymous):

25=5(4t-3) help please?

OpenStudy (potatoes.ramu):

25=20t-15 40=20t t=40/20 t=2

OpenStudy (texaschic101):

25 = 5(4t - 3) --- distribute through the parenthesis 25 = 20t - 15 -- add 15 to both sides 25 + 15 = 20t 40 = 20t -- divide by 20 40/20 = t 2 = t Potatoes is correct :)

OpenStudy (anonymous):

She's always is.. :)

OpenStudy (potatoes.ramu):

Haha not always... :) But thank you @texaschic101 and @Joeyhilliv :)

OpenStudy (texaschic101):

no problem :)